Suppose there are two dice: a blue with 6, 6, 3, 3, 1, 1, respectively, on its faces, and a green with 5, 4, 4, 4, 2, 1. When the two die are tossed, the dice showing the largest number wins.
A) which die would you expect to be the winner in the long run? Explain
B) in 36,000 tosses, about how many times would you expect the blue die to defeat the green die? Explain.

Answer by solver91311(24713)



Expected value is the same for both.
2/6 times 6 = 12/6
2/6 times 3 = 6/6
2/6 times 2 = 2/6      Total 20/6

1/6 times 5 = 5/6
3/6 times 4 = 12/6
1/6 times 2 = 2/6
1/6 times 2 = 1/6      Total 20/6


Dead heat; it is a coin toss which will win on any given roll.

John

My calculator said it, I believe it, that settles it